A car traveling at 20m/s, convert the speed from m/s to km/h
Answer #1

Speed of car in m/s is=20 m/s

As

1 km=1000 m

So

1m=(1/1000)km

And

1 h=3600 seconds

So

1s=(1/3600)h

Then

Speed of car in km/h is=20 m/s=(20)[(1/1000)/(1/3600)]km=(203600)/(1000)=72 km/h